OpenStudy (anonymous):

3y+45=10y+17 solve for y

OpenStudy (anonymous):

Use mathway.com

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

y=4

OpenStudy (anonymous):

It gives you the answer

OpenStudy (anonymous):

get y by itself by -3y from both sides. Then subtract 17 from both sides, that gives you 7y=28 then divide both sides by 7 and you get 4.

OpenStudy (anonymous):

3y+45=10y+17 45-17=10y-3y 28 = 7y y = 28/7 therefore y = 4

OpenStudy (anonymous):

well you just go through the numbers for y such as y=1, y=2, y=3 and work out the probem untill you get your answer

OpenStudy (anonymous):

FULL EXPLANATION AND ANSWER 3y+45=10y+17 Since 10y contains the variable to solve for, move it to the left-hand side of the equation by subtracting 10y from both sides. 3y+45-10y=17 Since 3y and -10y are like terms, add -10y to 3y to get -7y. -7y+45=17 Since 45 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 45 from both sides. -7y=-45+17 Add 17 to -45 to get -28. -7y=-28 Divide each term in the equation by -7. -(7y)/(-7)=-(28)/(-7) Simplify the left-hand side of the equation by canceling the common factors. y=-(28)/(-7) Simplify the right-hand side of the equation by simplifying each term. y=4